OpenStudy (anonymous):

Factor d^2 - 12d + 32 (d - 4) (d + 8) right?

OpenStudy (anonymous):

d^2 - 12d + 32 (d - 4) (d + 8) right?

OpenStudy (anonymous):

- in both brackets

OpenStudy (anonymous):

Could you explain how?

OpenStudy (acxbox22):

FOIL method

OpenStudy (lucaz):

it should be both negative because the sum of them must be equal to -12

OpenStudy (whpalmer4):

\[d^2-12d+32\]\[(d+a)(d+b) = d(d+b) + a(d+b) = d^2 + bd + ad + ab\]\[\qquad=d^2+ (a+b)d + ab\] Let's compare: \[d^2-12d+32\]\[d^2+(a+b)d + ab\] Those will be equal if we choose \(a,b\) such that \(a+b=-12\) and \(ab = 32\) So, we need a pair of factors of \(32\) that sum to \(-12\). \(-8,-4\) is such a pair:\[-8*-4 = 32\]\[-8-4 = -12\] Our factoring is therefore\[(d-4)(d-8) = d^2-8d-4d+32 = d^2-12d+32\]